Q4. Assume that the distance that a car runs on one liter of petrol varies inversely as the square of the speed at which it is driven. It gives a run of 25km per liter at a speed of 36km/h.At what speed should it be driven to get a run of 36km per liter ?
a. 34km/h
b. 33km/h
c. 31km/h
d. 30 km/h
1
Expert's answer
2013-03-15T08:48:42-0400
If a car gives a run of 25km per liter at a speed of 36km/h, as the distance that a car runs on one liter of petrol varies inversely as the square of the speed at which it is driven, then there exists constant C that

25/36² = C.

Therefore,

25/36² = C = 36/V²,

where V is the speed a car should be driven to get a run of 36km per liter. So,

V = √(36³/25) = 43.2.

There is no correct answer among the proposed ones. Correct answer is 43.2 km/h.